The wave function that describes a certain transverse wave at initial time t=0 is

y(x,t)=(7.20mm)sin2π(x/0.3m - t/0.0400s)

Find the wave's

a) amplitude

b) wavelength

c) frequency

d) speed of propagation

e) period

f) wave number

g) direction of propagation


1
Expert's answer
2021-01-29T04:29:15-0500

The wave function that describes a certain transverse wave 

"y(x,t)=y_{\rm max}\sin\left(\frac{2\pi}{\lambda}x-\frac{2\pi}{T}t\right)"

In our case

"y(x,t)=(7.20 \:{\rm mm}) \sin\left(\frac{2\pi}{0.3\:\rm m}x-\frac{2\pi}{0.04\:\rm s}t\right)"

a) amplitude

"y_{\rm max}=7.20 \:{\rm mm}"

b) wavelength

"\lambda=0.3\:\rm m"

c) frequency

"f=\frac{1}{T}=\frac{1}{0.04\:\rm s}=25\:\rm Hz"

d) speed of propagation

"v=f\lambda=25*0.3=7.5\:\rm m/s"

e) period

"T=0.04\:\rm s"

f) wave number

"k=\frac{2\pi}{0.3\:\rm m}=21\:\rm m^{-1}"

g) direction of propagation of the wave is negative "x" direction.